# Let Androids Dream of Electric Sheep: A Human-like Image Implication Understanding and Reasoning Framework

### v1

@@ 4.3.1 Implementation Details
MLLM-based Automatic Evaluation. To evaluate image implication comprehension in MLLMs, we develop an MLLM-based evaluation standard based on evaluation metrics, as illustrated in Figure 4. Our experiment utilize the same dataset from MCQ experiment, comprising 50 English images and 50 Chinese images. We employ human-written descriptions and implication interpretations as ground truth. We choose the same MLLMs with MCQ experiment to generate image implications for these images, which are subsequently scored using GPT-4o and our evaluation standard. The evaluation prompt is in Appendix D. To validate the model’s scoring efficacy, we enlist 16 PhD students and researchers well-versed in English and Chinese metaphorical imagery to independently score the dataset. The human-model scoring consistency reached 95.7%, affirming the method’s validity. The detailed human-model consistency study is in Appendix C.
@@ 4.3.2 Results and Analysis
Table 1 presents comprehensive results of OSQ across different MLLMs on our high-level benchmark. The LAD framework demonstrates exceptional effectiveness, achieving SOTA performance with the lightweight GPT-4o-mini model. In English OSQ, our framework substantially outperforms closed-sourced models like GPT-4o by 36.7% and Claude-3.5-Sonnet by 24.8%. For Chinese OSQ, while slightly below top closed-sourced models like GPT-4.1 and Doubao-1.5-thinking-vision-pro, our method still significantly surpasses Gemini-2.0-pro by 15.1% and DeepSeek-VL2 by 30%.
The enhancement over the GPT-4o-mini is particularly noteworthy, with improvements of 34.9% for English and 8.9% for Chinese, far exceeding other open-source and reasoning models. Unlike MCQ results, we observe significant performance disparities between reasoning and general models on OSQ, highlighting the distinct challenges of image implication generation. Interestingly, several models (e.g., QwenVL-2.5-72B, Gemini-2.0-pro) exhibit substantial performance gaps between MCQ and OSQ. Upon manual examination of model outputs, we attribute this to potential overfitting to multiple-choice formats and insufficient exposure to open-style generation tasks. In addition, LLMs or even MLLMs may not genuinely understand the questions but rather predict options as answers, introducing evaluation bias and demonstrating sensitivity to option positioning [38].
@@ 4.4.1 Stage I (Perception) and Stage III (Reasoning)
Table 2: Results of different CoT methods. Our LAD-CoT method achieves the best improvement. The best-improvement method in each category is in-bold.
@@
Compared to standard CoT, the results are shown in Table 2. While standard CoT offers minor gains in English (MCQ: 44% to 50%; OSQ: 2.98 to 3.10), it shows no improvement or even a slight decline in Chinese (MCQ: 42% unchanged; OSQ: 3.36 to 3.28). In contrast, LAD-CoT substantially outperforms both the baseline and standard CoT across all types. For instance, LAD-CoT achieves 68% on English MCQ while standard CoT only 50%, and a score of 3.84 on English OSQ compared to 3.10 for standard CoT. These findings highlight the superior efficacy of our LAD-CoT for image implication over standard CoT methods. A case study of various CoT on MCQ is in Figure 3. The standard CoT prompt and other details is in Appendix D.
@@ 4.4.2 Stage II (Search)
Grok-Search 72$\uparrow$ 64% 3.25 $\uparrow$ 2.92 $\downarrow$
@@
Perplexity (pro) 80$\uparrow$ 66$\uparrow$ 2.88 $\downarrow$ 3.28 $\downarrow$
@@
Table 3: Results of different search methods. Our LAD-Search method achieves the best improvement. The best-improvement method in each category is in-bold.
@@
Compared with Grok-3-search [30], GPT-4o-mini-search-preview, and GPT-4o with Perplexity.ai (Pro version), the results are shown in Table 3. GPT-Search, when applied to GPT-4o-mini, improves MCQ scores but degrades OSQ performance (English OSQ: 3.84 to 3.62, Chinese OSQ: 3.58 to 3.34). Grok-Search, on the Grok-3 model, provides limited gains, mainly in English MCQ (66% to 72%), exhibits inconsistent Chinese performance, and shows minimal OSQ improvement. Perplexity.ai search with GPT-4o significantly boosts MCQ accuracy, but it markedly lowers OSQ scores (English OSQ: 2.94 to 2.88, Chinese OSQ: 3.76 to 3.28). In contrast, LAD-Search consistently enhances performance across both MCQ and the more challenging OSQ. This underscores its superior ability to effectively integrate external knowledge for implication understanding, outperforming other search methods particularly in open-style reasoning scenarios where they often falter.
@@ 5.1 How to Let Androids Dream? Perception and Reasoning
The question “How to Let Androids Dream?” metaphorically addresses the foundational challenge of enabling AI systems to interpret the nuanced implications embedded in images. Our framework tackles this by first emulating human-like perception (Stage I), converting raw visual input into rich, multi-level textual representations, including comprehensive descriptions and salient keywords. These keywords are designed to capture not only objects and scenes but also potential emotional tones, relevant domains (e.g., cultural, social, political), and discernible rhetorical devices. Subsequently, LAD’s Stage III employs an explicit, structured CoT process. This structured reasoning guides the model to systematically connect the perceived visual elements with retrieved contextual knowledge, thereby constructing a coherent understanding of implications. This method is vital because, as our experiments (Section 4) and recent work on social reasoning [10] show, comprehending implications extends beyond basic VQA tasks and classic logical reasoning; it inherently involves sophisticated social reasoning and the interpretation of contextual cues often missed by MLLMs.
@@ 6 Conclusion
Understanding image implications remains challenging for MLLMs, mainly due to missing contextual information. Our work introduces LAD, a novel three-stage framework—Perception, Search, and Reasoning. Inspired by human cognitive processes, this framework is designed to achieve contextual alignment by explicitly integrating visual interpretation with external knowledge retrieval. We conduct comprehensive experiments to demonstrate its effectiveness. Utilizing the lightweight GPT-4o-mini, LAD achieves SOTA results on English and Chinese implication benchmarks, performing comparable or even surpassing GPT-4o and other top closed-source models, particularly on challenging Open-Style Question. In summary, LAD bridges the gap between superficial perception and deep reasoning in multimodal AI systems, offering a promising direction for context-aware reasoning.
@@
1) The search stage, particularly the websearch and multiple model calls, will make latency in generating image implications.
@@ Appendix A Algorithm
Input: Image $IMG$, Task $T_{MCQ}$, Task $T_{OSQ}$
@@ Appendix B Experiment Setup
We use the lightweight GPT-4o-mini-0718 [19] with LAD framework in experiments. We set the model temperature as 0.5 and top_p as 0.9 in MCQ experiments, and temperature as 0.7 and top_p as 0.9 in OSQ experiments. Additionally, we set the evaluation model GPT-4o temperature as 0 and evaluate more than three times to get the average score in OSQ experiments. All experiments are conducted on NVIDIA A800 GPUs.
@@ Appendix C Human-Model Consistency Study
To validate our automated OSQ evaluation based on the GPT-4o model, we conduct a human-model consistency study. We construct a dedicated dataset by randomly selecting 25 images with questions each from our English and Chinese OSQ. We recruit 16 PhD students and researchers, all proficient in both English and Chinese and experienced with metaphorical imagery, to independently score the model responses. Their evaluations are based on ground truth answers and detailed scoring standard. We calculate human inter-annotator agreement by averaging the scores for each response after discarding the highest and lowest individual scores. This process yields the consistency of 94.8% for Chinese and 96.5% for English. The average human-model scoring consistency reached 95.7%, affirming the method’s validity for assessing image implication comprehension.
Appendix D Prompts

### v2

@@ 4.3.1 Implementation Details
MLLM-based Automatic Evaluation. To evaluate image implication comprehension in MLLMs, we develop an MLLM-based evaluation standard based on evaluation metrics, as illustrated in Figure 4. Our experiment utilize the same dataset from MCQ experiment, comprising 50 English images and 50 Chinese images. We employ human-written descriptions and implication interpretations as ground truth. We choose the same MLLMs with MCQ experiment to generate image implications for these images, which are subsequently scored using GPT-4o and our evaluation standard. The evaluation prompt is in Appendix F. To validate the model’s scoring efficacy, we enlist 16 PhD students and researchers well-versed in English and Chinese metaphorical imagery to independently score the dataset. The human-model scoring consistency reached 95.7%, affirming the method’s validity. The detailed human-model consistency study is in Appendix C.
@@ 4.3.2 Results and Analysis
Table 1 presents comprehensive results of OSQ across different MLLMs on our high-level benchmark. The LAD framework demonstrates exceptional effectiveness, achieving SOTA performance with the lightweight GPT-4o-mini model. In English OSQ, our framework substantially outperforms top closed-sourced models like Gemini-3.0-pro by 5.2%, GPT-4o by 36.7% and Claude-3.5-Sonnet by 24.8%. For Chinese OSQ, while slightly below top closed-sourced models like Gemini-3.0-pro and Doubao-1.5-thinking-vision-pro, our method still significantly surpasses Qwen2.5-VL-72B by 15.1% and DeepSeek-VL2 by 30%.
The enhancement over the GPT-4o-mini is particularly noteworthy, with improvements of 34.9% for English and 8.9% for Chinese, far exceeding other open-source and reasoning models. Unlike MCQ results, we observe significant performance disparities between reasoning and general models on OSQ, highlighting the distinct challenges of image implication generation. Interestingly, several models (e.g., Qwen2.5-VL-72B, Gemini-2.0-flash) exhibit substantial performance gaps between MCQ and OSQ. Upon manual examination of model outputs, we attribute this to potential overfitting to multiple-choice formats and insufficient exposure to open-style generation tasks. In addition, LLMs or even MLLMs may not genuinely understand the questions but rather predict options as answers, introducing evaluation bias and demonstrating sensitivity to option positioning [zheng2024largelanguagemodelsrobust].
@@ 4.4.1 Stage I (Perception) and Stage III (Reasoning)
Table 2: Results of different CoT methods. Our LAD-CoT method achieves the best improvement. The best-improvement method in each category is in-bold. Performance differences relative to base models are shown as colodarkred subscripts: for improvements, for declines.
@@
Compared to standard CoT, the results are shown in Table 2. While standard CoT offers minor gains in English (MCQ: 44% to 50%; OSQ: 2.98 to 3.10), it shows no improvement or even a slight decline in Chinese (MCQ: 42% unchanged; OSQ: 3.36 to 3.28). In contrast, LAD-CoT substantially outperforms both the baseline and standard CoT across all types. For instance, LAD-CoT achieves 68% on English MCQ while standard CoT only 50%, and a score of 3.84 on English OSQ compared to 3.10 for standard CoT. These findings highlight the superior efficacy of our LAD-CoT for image implication over standard CoT methods. A case study of various CoT on MCQ is in Figure 3. The standard CoT prompt and other details is in Appendix F.
@@ 4.4.2 Stage II (Search)
Grok-Search 72% $\uparrow$ 64% 3.25 $\uparrow$ 2.92 $\downarrow$
@@
Perplexity (pro) 80% $\uparrow$ 66% $\uparrow$ 2.88 $\downarrow$ 3.28 $\downarrow$
@@
Table 3: Results of different search methods. Our LAD-Search method achieves the best improvement. The best-improvement method in each category is in-bold. Performance differences relative to base models are shown as colodarkred subscripts: for improvements, for declines.
@@
Compared with Grok-3-search [grok3], GPT-4o-mini-search-preview, and GPT-4o with Perplexity.ai (Pro version), the results are shown in Table 3. GPT-Search, when applied to GPT-4o-mini, improves MCQ scores but degrades OSQ performance (English OSQ: 3.84 to 3.62, Chinese OSQ: 3.58 to 3.34). Grok-Search, on the Grok-3 model, provides limited gains, mainly in English MCQ (66% to 72%), exhibits inconsistent Chinese performance, and shows minimal OSQ improvement. Perplexity.ai search with GPT-4o significantly boosts MCQ accuracy, but it markedly lowers OSQ scores (English OSQ: 2.94 to 2.88, Chinese OSQ: 3.76 to 3.28). In contrast, LAD-Search consistently enhances performance across both MCQ and the more challenging OSQ. This underscores its superior ability to effectively integrate external knowledge for implication understanding, outperforming other search methods particularly in open-style reasoning scenarios where they often falter.
4.4.3 Different Base Models
Model Multiple-Choice Question Open-Style Question
en zh en zh
Qwen2.5-VL-7B
w/o LAD 46% 40% 2.34 2.58
w/ LAD 64% $\uparrow$ 46% $\uparrow$ 3.64 $\uparrow$ 3.36 $\uparrow$
Qwen2.5-VL-72B
w/o LAD 72% 56% 1.56 3.12
w/ LAD 76% $\uparrow$ 62% $\uparrow$ 3.62 $\uparrow$ 3.68 $\uparrow$
GPT-4o
w/o LAD 74% 58% 2.94 3.76
w/ LAD 80% $\uparrow$ 66% $\uparrow$ 4.14 $\uparrow$ 4.26 $\uparrow$
Gemini-3.0-pro
w/o LAD 76% 76% 3.82 3.96
w/ LAD 82% $\uparrow$ 78% $\uparrow$ 4.30 $\uparrow$ 4.46 $\uparrow$
Table 4: Results of different base models. Our LAD demonstrates the generalizability on different base models. The best-performing model in each category is in-bold. Performance differences relative to base models are shown as colodarkred subscripts: for improvements, for declines.
To demonstrate the generalizability of our LAD framework beyond the GPT-4o-mini model, we conduct new experiments applying LAD to other base models, including the open-source Qwen2.5VL series, the closed-source model GPT-4o and the latest closed-source model Gemini-3.0-pro. As the Table 4 shows, applying LAD framework significantly boosts the performance of all models across both MCQ and OSQ tasks, confirming that our framework is not model-specific and provides a robust and generalizable approach to enhancing image implication understanding.
4.4.4 Generalization Experiment
Model Multiple-Choice Question Open-Style Question
II-Bench (1399) CII-Bench (800) II-Bench (1399) CII-Bench (800)
GLM-4.1V-8B 70.0% 46.3% 2.83 3.06
GPT-4o-mini 63.5% 35.6% 2.93 3.29
InternVL3-78B 78.2% 64.0% 3.68 4.06
GPT-4o 72.6% 54.1% 3.86 4.06
Claude-3.5-Sonnet 80.9% 54.1% 3.51 3.84
LAD (GPT-4o-mini) 81.2% $\uparrow$ 53.8% $\uparrow$ 4.22 $\uparrow$ 4.31 $\uparrow$
Table 5: Results of different models on full benchmarks. The best-performing model in each category is in-bold. Performance differences relative to base models are shown as colodarkred subscripts: for improvements, for declines.
Model MMMU_val SeedBench MMStar
GPT-4o-mini 59.4 72.8 54.8
GPT-4o 70.7 76.7 65.1
LAD (GPT-4o-mini) 67.9 % $\uparrow$ 77.2% $\uparrow$ 60.3 $\uparrow$
Table 6: Results of different models on general VQA benchmarks. The best-performing model in each category is in-bold. Performance differences relative to base models are shown as colodarkred subscripts: for improvements, for declines.
Experiments On Full Benchmarks. We conduct the large-scale experiments with the representative and top-performing models, including Closed-Source models GPT-4o and Claude-3.5-Sonnet, as well as the Open-Source model GLM-4.1V-8B, on the full benchmarks: II-Bench (1,399 examples) and CII-Bench (800 examples) for both MCQ and OSQ tasks.
As the results in Table 5 show, our LAD framework’s significant performance gains are consistent on these much larger datasets. Notably, by applying LAD, the lightweight GPT-4o-mini significantly surpasses the much larger GPT-4o and Claude-3.5-Sonnet. Compared with the baseline GPT-4o-mini model, we can find that: (1) On the large-scale English benchmark (II-Bench), our LAD framework improves the GPT-4o-mini score from 63.5% to 81.2% on MCQ and 2.93 to 4.22 on OSQ. This is a substantial absolute increase of 17.7% (27.9% relative improvement) and 1.29 (44% relative improvement). (2) The gains on the large-scale Chinese benchmark (CII-Bench) are even more pronounced. LAD boosts performance from 35.6% to 53.8% on MCQ and 3.29 to 4.31 on OSQ, representing an absolute increase of 18.2% (51.1% relative improvement) and 1.02 (31% relative improvement).
This robust improvement is consistent with the trend we observed and reported on our high-level benchmark (smaller 100-image dataset) in Table 1. While the exact percentages differ due to the varying scales and baselines of the datasets, the key takeaway is that the significant positive impact of the LAD framework is undeniable across both small and large-scale evaluations. This analysis confirms that our framework’s benefits are not an artifact of a small test set but are indeed robust and generalizable. It also reflects the reliability and high quality of our manually curated high-level benchmark.
Experiments On General VQA Benchmarks. To further demonstrate that LAD is a generalizable reasoning framework, we evaluated it on three general multi-modal benchmarks: MMMU (Expert AGI and Visual Reasoning), SeedBench (General Understanding), and MMStar (General Understanding). We applied the LAD framework to GPT-4o-mini without modifying the core architecture. The results are presented in Table 6.
We find that the LAD framework provides huge improvements (e.g., +8.5% on MMMU). With LAD, the lightweight GPT-4o-mini surpasses the much larger GPT-4o on SeedBench (77.2 vs 76.7) and significantly closes the gap on others. These results confirm that our "Perception-Search-Reasoning" workflow addresses a fundamental cognitive gap in VLM reasoning, effectively handling tasks requiring visual commonsense and complex reasoning beyond just metaphor understanding.
@@ 5 Discussion
5.1 Human Cognitive Theory of Let Androids Dream
Our claim is that the LAD framework is analogous to human cognitive strategies, not a direct neuroscientific replica. Our goal is to create a system that reasons in a way that is transparent and aligns with how humans might tackle the same problem, not to simulate the human brain perfectly.
Our framework is directly inspired by established human cognitive science theories: (1) Dual-Process Theory [Evans2003InTM]: The Perception stage mirrors the interplay between System 1 (the fast, intuitive, holistic impression of the image) and System 2 (the slower, analytical identification of key elements), and (2) Active Information-Seeking Theory [Informationseeking, Wilson2009ActivityTA]: The Search stage is analogous to the human tendency to actively seek external information to resolve ambiguity. Humans do not reason in a vacuum; when we encounter an unfamiliar meme or cultural reference, a common cognitive act is to "Google it" to supplement our internal knowledge. Our WebSearch module directly simulates this deliberate information-foraging behavior.
@@ 5.2 How to Let Androids Dream? Perception and Reasoning
The question “How to Let Androids Dream?” metaphorically addresses the foundational challenge of enabling AI systems to interpret the nuanced implications embedded in images. Our framework tackles this by first emulating human-like perception (Stage I), converting raw visual input into rich, multi-level textual representations, including comprehensive descriptions and salient keywords. These keywords are designed to capture not only objects and scenes but also potential emotional tones, relevant domains (e.g., cultural, social, political), and discernible rhetorical devices. Subsequently, LAD’s Stage III employs an explicit, structured CoT process. This structured reasoning guides the model to systematically connect the perceived visual elements with retrieved contextual knowledge, thereby constructing a coherent understanding of implications. This method is vital because, as our experiments (Section 4) and recent work on social reasoning [kim2025hypothesisdriventheoryofmindreasoninglarge] show, comprehending implications extends beyond basic VQA tasks and classic logical reasoning; it inherently involves sophisticated social reasoning and the interpretation of contextual cues often missed by MLLMs.
@@ 6 Conclusion
Understanding image implications remains challenging for MLLMs, mainly due to contextual missing. Our work introduces Let Androids Dream (LAD), a novel three-stage framework: Perception, Search, and Reasoning. Inspired by human cognitive processes, this framework is designed to achieve contextual alignment by explicitly integrating visual interpretation with external knowledge retrieval. We conduct comprehensive experiments to demonstrate its effectiveness. Utilizing the lightweight GPT-4o-mini, LAD achieves top results on implication benchmarks, performing comparable or even surpassing Gemini-3.0-pro and other top closed-source models, particularly on challenging OSQ. In summary, LAD bridges the gap between superficial perception and reasoning in multimodal AI systems, offering a promising direction for contextual-alignment vision-language reasoning.
@@
1) The search stage, particularly the websearch and multiple model calls, will incur small latency in generating image implications. Based on our experiments, a single search question takes approximately 35s to 55s and the whole search stage takes 3 mins to 5 mins to process through the entire pipeline. The process consumes between 3,440 to 4,280 tokens per image.
@@ Appendix A Algorithm
Input: Image IMGIMG, Task $T_{MCQ}$, Task $T_{OSQ}$
@@ Appendix B Experiment Setup
We use the lightweight GPT-4o-mini-0718 [gpt4o] with LAD framework in experiments. We set the model temperature as 0.5 and top_p as 0.9 in MCQ experiments, and temperature as 0.7 and top_p as 0.9 in OSQ experiments. Additionally, we set the evaluation model GPT-4o temperature as 0 and evaluate more than three times to get the average score in OSQ experiments. All experiments are conducted on NVIDIA A800 GPUs.
@@ Appendix C Human-Model Consistency Study
To validate our automated OSQ evaluation based on the GPT-4o model, we conduct a human-model consistency study. We construct a dedicated dataset by randomly selecting 25 images with questions each from our English and Chinese OSQ. We recruit 16 PhD students and researchers, all proficient in both English and Chinese and experienced with metaphorical imagery, to independently score the model responses. Their evaluations are based on ground truth answers and the detailed scoring standard. We calculate human inter-annotator agreement by averaging the scores for each response after discarding the highest and lowest individual scores. This process yields the consistency of 94.8% for Chinese and 96.5% for English. The average human-model scoring consistency reached 95.7%, affirming the method’s validity for assessing image implication comprehension.
Appendix D Statistics
We manually construct the high-level benchmark by selecting 100 high-quality, diverse and representative images from II-Bench [liu2024iibenchimageimplicationunderstanding] and CII-Bench [zhang2024mllmsunderstanddeepimplication]. The general statistic is in Table 7.
Statistics of English Images
Society 21 (42%)
Life 16 (32%)
Art 6 (2%)
Psychology 4 (8%)
Others 3 (6%)
Multi-panel Comic 16 (32%)
Single-panel Comic 9 (18%)
Illustration 5 (10%)
Meme 5 (10%)
Poster 5 (10%)
Painting 5 (10%)
Logo 5 (10%)
Statistics of Chinese Images
Life 13 (26%)
Art 13 (26%)
Society 12 (24%)
Chinese Traditional Culture 6 (12%)
Environment 5 (10%)
Politics 1 (2%)
Illustration 15 (30%)
Single-panel Comic 10 (20%)
Poster 8 (16%)
Meme 8 (16%)
Painting 6 (12%)
Multi-panel Comic 3 (6%)
Table 7: General statistics of the high-level benchmark.
Appendix E Further Analysis on Method and Experiments
E.1 Analysis of Let Androids Dream Success
Our analysis points to two primary failure modes for baseline models, which Let Androids Dream (LAD) is designed to mitigate. These are illustrated in Figure 1 and the case study in Figure 3:
1. Superficial Reasoning: This occurs when a model only processes the literal, surface-level elements and misses the metaphorical meaning entirely. In Figure 3 the "End2End" baseline exemplifies this, failing to grasp the subversion of the fairy tale trope.
2. Over-Inference: This happens when a model incorrectly applies a known symbol or narrative without considering the full context. The “CoT" baseline in Figure 3 demonstrates this by connecting the heart symbol to a traditional fairy tale transformation without recognizing the comic’s twist.
LAD succeeds by first creating a more structured understanding in the Perception stage and then grounding its reasoning with targeted external knowledge from the Search stage, which helps avoid both superficiality and incorrect inferences.
E.2 Analysis of Model Scaling and Image Implication Types
Our experiments have some insightful findings:
1. Model Scaling: By testing on QwenVL-2.5-7B and QwenVL-2.5-72B, we can analyze the effect of model scale. Our findings align with expectations: larger parameter models generally achieve better baseline performance, and both scales benefit from the LAD framework. This confirms that our method is effective across different model sizes.
2. Image Implication Types: Our benchmark was already designed to be diverse across various domains (e.g., life, society, art, psychology, Chinese traditional culture) and image types (e.g., comic, poster, meme). We find that models perform worse in domains containing abstract and complex information, like Art and Psychology. And models only observe the surface-level information and lack sufficient understanding of Chinese culture. In a further analysis using the annotations from the original II-Bench and CII-Bench, we observed that providing explicit labels for Emotion, Domain, and Rhetoric significantly enhances model accuracy, with Emotion labels providing the largest boost. This confirms that our framework’s focus on identifying these elements in the Perception stage is well-founded.
Appendix F Prompts